OpenStudy (anonymous):

9 3/4 percent converted to a fraction equals

OpenStudy (australopithecus):

9 + 3/4 9(4)/4 + 3/4

OpenStudy (australopithecus):

(9(4) + 3)/4

OpenStudy (diyadiya):

\[9 \frac{3}{4}= 9+\frac{3}{4}= \frac{9 \times 4 +3}{4}= \frac{36+3}{4}= \frac{39}{4}\]

OpenStudy (australopithecus):

remmeber that 3 = 3(4)/(4) = 3(1) = 3

OpenStudy (australopithecus):

you are just manipulating the expression without changing anything so that you can add the fractions

OpenStudy (australopithecus):

because remember to add or subtract fractions they must have the same denominators

OpenStudy (asnaseer):

if the question is to convert:\[9\frac{3}{4}\text{%}\]to a fraction, then you need to remember that percent means "out of 100".

OpenStudy (asnaseer):

so:\[9\frac{3}{4}\text{%}=\frac{39}{4}\div100=\frac{39}{400}\]

OpenStudy (asnaseer):

for example to convert 50% to a fraction you would do:\[50\text{%}=50\div100=\frac{50}{100}=\frac{1}{2}\]

OpenStudy (asnaseer):

ali322n: please ask questions if you need more explanation

OpenStudy (diyadiya):

@ali322n Sorry for the wrong answer.. didn't read the question properly Thanks @asnaseer

OpenStudy (asnaseer):

Diyadiya: np - we all make mistakes - that's what makes us human :)

OpenStudy (diyadiya):

:D

OpenStudy (anonymous):

okay thank you asnaseer

OpenStudy (asnaseer):

yw